Determine the number of different three letter "words" that can be formed using the letters from the word "ARRANGE": Determine the number of different three letter "words" that can be formed using the letters from the word "ARRANGE"
Answer by edjones(2415)

1st use each letter only once in the 3 letter word:
5P3=5!/(5-3)!=(5*4*3*2)/2=60
2nd use both double letters with one other letter. rrn rnr nrr There are 3 permutations for each of the 4 single letters=12 There are 2 sets of double letters so 24 are the possible permutations.
60+24=84 "words" altogether.
.
Ed